# Analysis of Internet of Things implementation barriers in the cold supply chain: an integrated ISM-MICMAC and DEMATEL approach

### v2

@@ Abstract
Integrating Internet of Things (IoT) technology inside the cold supply chain can enhance transparency, efficiency, and quality, optimizing operating procedures and increasing productivity. The integration of IoT in this complicated setting is hindered by specific barriers that need a thorough examination. Prominent barriers to IoT implementation in the cold supply chain are identified using a two-stage model. After reviewing the available literature on the topic of IoT implementation, a total of 13 barriers were found. The survey data was cross-validated for quality, and Cronbach’s alpha test was employed to ensure validity. This research applies the interpretative structural modeling technique in the first phase to identify the main barriers. Among those barriers, “regularity compliance” and “cold chain networks” are key drivers for IoT adoption strategies. MICMAC’s driving and dependence power element categorization helps evaluate the barrier interactions. In the second phase of this research, a decision-making trial and evaluation laboratory methodology was employed to identify causal relationships between barriers and evaluate them according to their relative importance. Each cause is a potential drive, and if its efficiency can be enhanced, the system as a whole benefits. The research findings provide industry stakeholders, governments, and organizations with significant drivers of IoT adoption to overcome these barriers and optimize the utilization of IoT technology to improve the effectiveness and reliability of the cold supply chain.
@@ 1 Introduction
The supply chain (SC) is a network of related entities that includes every procedure step, starting with the procurement of raw materials or components to the consummation of the finished good. The cold SC (CSC) is a procedure that includes activities regulated by temperature and a specialized logistical infrastructure designed to maintain product quality and temperature consistency throughout its entire transit, beginning with the point of origin in manufacturing or procurement and ending with its eventual delivery to the final consumer [1]. Everything from preparing, storing, transporting, and monitoring temperature-sensitive goods from their providing areas to their receiving locations involves a sequence of activities in the CSC. Food, flowers, medicines, and chemicals are all examples of temperature-sensitive items because of their perishable nature [2].
@@
The Internet of Things (IoT) refers to a network that links physical items to the Internet and allows them to broadcast data in order to intelligently identify themselves and enable their location, tracking, monitoring, and management with the assistance of radio frequency identification tags, sensors, actuators, and positioning systems [3]. The IoT emerged from networked devices, which started at MIT’s Auto-ID Centre in the nineties, according to Centre Director Kevin Ashton, who purportedly developed the concept of IoT at the turn of the millennium. The current concept of IoT is backed by the addition of GPS devices, smartphones, social networks, cloud computing, and data analytics to these existing networks [4]. The IoT plays a significant part in Industry 4.0, which encompasses a variety of innovative and revolutionary technologies, cyber-physical systems, the IoT, and cloud computing—in particular [5]. It links objects, sensors, actuators, and other smart technology, allowing immediate access to physical commodities data to support new services with great efficiency and productivity [6].
@@
Improved operating processes, reduced risk and expense, more visibility and transparency, and greater adaptability and flexibility across the supply chain are merely a few of the many advantages of incorporating IoT into the supply chain [7]. IoT technology may drastically boost the monitoring and control of CSCs, particularly in areas like medications, food, and logistics where temperature-sensitive commodities are moved and preserved [8].
@@
Integrated Interpretive Structural Modeling and Decision Making Trial and Evaluation Laboratory (ISM-DEMATEL) is an extensively used technique for decision-making and analysis in complex systems. ISM is beneficial for building hierarchical links among factors and analyzing the structure of complex systems, while DEMATEL delivers a quantitative assessment of the intensity and direction of interactions between factors [9][10]. Both approaches allow for a combination of qualitative and quantitative research, with DEMATEL specifically addressing feedback loops. They are straightforward and obvious, making them easier to explain and grasp by decision-makers [11][12][13].
@@
The increasing volume of international trade and consumer spending has spurred the fast growth of the supply chains for perishable commodities. Logistics for “cold chain” items, such as perishables and frozen foods, constantly evolve. Global revenue for the sector is forecast to reach USD 277.69 billion by 2023, growing at a CAGR (cumulative aggregate gross revenue) of 7.24 percent between 2018 and 2023 [14].
@@
A simplified and technology-enhanced logistics method for CSC IoT implementation is shown in Fig 1. The graphic depicts IoT sensors installed across the CSC to track moisture, temperature, and other important factors, including refrigerated vehicles, warehouses, and storage facilities. To provide immediate insight into the state of the products in transit, data collected from these sensors is wirelessly connected and relayed in real-time to a centralized system [15]. In sectors like pharmaceuticals and food, where maintaining the proper temperature is crucial to product quality and safety, IoT devices make sure that temperature-sensitive commodities are stored and transported within the designated temperature range [16]. By considering traffic, weather, and the quality of the supplies, IoT may help enhance delivery routes [17]. This guarantees that perishable goods arrive on schedule and in the best possible shape. IoT sensors can monitor inventory levels to minimize overstocking or understocking of temperature-sensitive products, which lowers spoilage and perishable goods losses [18]. Well-organized logistics and transportation systems greatly affect product quality. Adopting a high-quality logistics system, especially for perishable items, involves frequent temperature monitoring, real-time product quality assessment, and shelf life determination [19].
Several studies have delved into barriers within specific countries and contexts; however, these investigations are not without limitations. Some have explored sectors unrelated to the CSC, and while certain studies have focused on IoT adoption in this context, they often fall short in establishing connections among identified barriers or evaluating their respective impacts on the overall system. Moreover, the exploration of the IoT implementation barriers in CSCs represented a significant gap in the existing literature, with no prior comprehensive studies identified in this specific domain. The lack of scholarly attention to this crucial aspect within CSC underscored the need for an in-depth investigation. Our research stands as a pioneering effort to bridge this notable gap by employing an integrated ISM-DMATEL approach.
The literature gap is particularly evident in the scarcity of empirical examinations and theoretical frameworks dedicated to understanding the intricacies of implementing IoT in CSCs. The unique challenges posed by the CSC environment, characterized by stringent temperature control requirements and perishable goods logistics, warranted a dedicated exploration.
To address these gaps, this paper contributes to the existing literature in the following ways:
@@
    In-depth analysis of research on essential barriers to CSC IoT adoption, considering various perspectives, including supply chain stages, environmental effects, socio-economic scenery, and data complexity.
@@
    Employing a survey-based approach to gather data and expert opinions on these barriers.
@@
    The application of the MICMAC strategy to comprehensively analyze the fundamental constructs, thereby offering valuable insights into their significance.
@@
    The utilization of the DEMATEL approach to validate the causal interrelationships among the components within the model.
@@
    Highlighting key barriers that play a fundamental role in efficiently managing the cold chain, offering insights to overcome challenges in IoT adoption.
@@
    Systematic identification and ranking of barriers based on their influence on the system, addressing the current lack of comprehensive assessments in the literature.
@@
    Providing a structured approach to understanding and addressing the challenges associated with IoT implementation in the CSC, contributing to enhanced operational efficiency and reliability.
@@
    Offering valuable insights for enterprises, governments, and technology developers to navigate and overcome obstacles in integrating IoT into CSC operations.
The rest of the paper is arranged as follows: “Literature review” provides a comprehensive literature review of IoT implementation in CSCs, including implementation challenges and comparable methodologies utilized in the present era; “Methodology” presents the research methodology that was used for the study; “Numerical illustration for results interpretation” covers the data analysis, along with the application of ISM-DEMATEL to the obtained barriers; “Discussions” describes the result of the implementation of ISM-DEMATEL techniques and discussion about the result. Finally, the study finishes with conclusions, limitations, and recommendations for future work in “Conclusions and future works”.
@@ 2 Literature review
This is part of the article’s aim to provide a summary of how cold SCM procedures may make use of IoT. We have used Google Scholar and many publishers’ websites, including Springer, Elsevier, Emerald, and Taylor-and-Francis, to compile the literature most relevant to our study.
@@ 2.1 Internet of Things (IoT)
First invented in 1999 by Kevin Ashton and many other scholars at the Massachusetts Institute of Technology, the IoT has gained broad interest because it can provide comfort, efficiency, and competitive benefits to the corporate sector [20]. IoT is an essential idea in the coming period. Numerous smart buildings and smart houses are being built using IoT technologies [21]. IoT presents a new opportunity to reduce risks, manage complexity, and produce concrete economic benefits via increasing transparency and flexibility across the supply chain [22]. The physical components that link the IoT include wireless devices like portable computers, smartphones, wearable gadgets, RFID tags, RFID readers, wireless sensors, etc. Users may interact with the IoT network using a variety of devices, including cellphones, implanted RFID transponders, wirelessly connected PCs, and wearable tech [23]. The popularity and success of IoT have skyrocketed in recent years. Industry leaders, technologists, and consumers see the importance of IoT, which has led to an industrial revolution and automated operations in organizations and homes [24]. Human-to-thing communication and autonomous coordination among “things” is possible even when they are kept at a facility or transported from one entity in the supply chain to another. Increased supply chain visibility, responsiveness, and adaptability to a variety of SCM challenges are all made possible by the IoT [25] [4].
@@ 2.2 IoT in CSC
The IoT, often known as the third wave of IT, has revolutionized supply chain management (SCM) and other sectors. This helps the supply chain function expand by offering technical assistance for greater visibility, stability, and intelligent management [26]. Ben-Daya et al. [4] analyzed the existing literature to learn more about the IoT and how it influences SCM. In addition, the majority of research has concentrated on the steps in the supply chain that pertain to food production, distribution, and processing. Kamble et al. [27] proposed an integrated ISM- DEMATEL model to identify the most significant barriers to the widespread deployment of IoT technologies in the food retail supply chain. The two most significant barriers to the widespread use of the IoT were found to be a lack of government regulation and an insufficient network infrastructure.
@@
Numerous deployment strategies, constraints, and barriers in cold chain monitoring applications were discussed [28]. The feasibility of implementing modern technologies such as the IoT to enhance the cold SC was assessed in this article [29]. In order to enhance its service quality, the organization must invest additional funds in maintenance. Umamaheswari et al. [30] suggested smart cold storage that uses cutting-edge supply chain technology and the IoT to boost productivity and reduce wait times at all levels. Luo et al. [31] presented a smart monitoring system that stores data on defined servers to monitor cold chain items in real time while cutting costs, enhancing network capacity, and simplifying protocols. Cold chain transit is easy to track using this strategy. Afreen and Bajwa [32] implemented a system that utilizes the IoT to track and notify users of changes in cold storage conditions such as humidity, temperature, light, and gas concentration in real-time., alerting workers when these variables exceed safe thresholds. Goodarzian et al. [33] proposed a model for improving vaccine injection’s precision, swiftness, and fairness and implemented the inaugural responsive green-cold vaccine supply chain network in response to the COVID-19 pandemic. Intelligent control technology, a cellular network, wireless communication technology, temperature and humidity sensors, the IoT, and database technology all come together to form the cold chain management system[34]. Ceken and Abdurahman [35] presented an IoT-based cold chain logistics system that controls and monitors the cold chain’s ambient temperature in real-time and estimates perishable item shelf life to help all stakeholders make better decisions. The paper by [18] presented an IoT-based risk monitoring solution to handle product quality and safety concerns in the cold chain. Using this technology, the cold chain can be monitored in real-time, and risks assessed to ensure the safety of workers. Hernandez and Yamaura [36] suggested that IoT-enabled devices may reduce pharmaceutical waste from temperature changes. It also showed consistency in IoT implementation, including infrastructure, security, and device management. Due to perishable food and cross-regional transit, cold chain systems encounter several challenges in perishable food supply chains. This study of [37] addressed these difficulties by creating a perishable food supply chain model utilizing real-time IoT data and two supply hubs to increase performance. Future cold supply networks can be more resilient and sustainable if they use IoT technologies. This advancement is expected to result in less waste and decreased operational expenses, therefore benefiting both businesses and consumers [29] [38].

### v4

# Analysis of Internet of Things Implementation Barriers in the Cold Supply Chain: An Integrated ISM–MICMAC and DEMATEL Approach

@@ Abstract
Integrating Internet of Things (IoT) technology inside the cold supply chain can enhance transparency, efficiency, and quality, optimize operating procedures, and increase productivity. The integration of the IoT in this complicated setting is hindered by specific barriers that require thorough examination. Prominent barriers to IoT implementation in a cold supply chain, which is the main objective, are identified using a two–stage model. After reviewing the available literature on IoT implementation, 13 barriers were identified. The survey data were cross–validated for quality, and Cronbach’s alpha test was employed to ensure validity. This study applies the interpretative structural modeling technique in the first phase to identify the main barriers. Among these barriers, “regulatory compliance” and “cold chain networks” are the key drivers of IoT adoption strategies. MICMAC’s driving and dependence power element categorization helps evaluate barrier interactions. In the second phase of this study, a decision-making trial and evaluation laboratory methodology was employed to identify causal relationships between barriers and evaluate them according to their relative importance. Each cause is a potential drive, and if its efficiency can be enhanced, the system benefits as a whole. The findings provide industry stakeholders, governments, and organizations with significant drivers of IoT adoption to overcome these barriers and optimize the utilization of IoT technology to improve the effectiveness and reliability of the cold supply chain.
@@ 1 Introduction
Recent developments in Internet of Things (IoT) technology have the potential to significantly change several industries via the introduction of data-driven insights, increased connectivity, and a replacement of long-established ways of doing things. The supply chain (SC) is a network of related entities that includes every procedure step, from the procurement of raw materials or components to the consummation of the finished good. Supply chains in the modern day have developed into intricate value networks and are now an important differentiator for businesses. A growing number of barriers make it harder than ever to track items and goods as they go through the value chain and confirm where they came from [1]. Cold SC (CSC) is a procedure that includes activities regulated by temperature and a specialized logistical infrastructure designed to maintain product quality and temperature consistency throughout its entire transit, beginning with the point of origin in manufacturing or procurement and ending with its eventual delivery to the final consumer [2]. Everything from preparing, storing, transporting, and monitoring temperature-sensitive goods from their providing areas to their receiving locations involves a sequence of activities in CSC. Food, flowers, medicines, and chemicals are examples of temperature-sensitive items because of their perishable nature [3]. To increase sales and profits, companies are creating a cold chain in transportation by reducing costs, reducing delivery times, improving product quality, and offering personalized goods.
@@
The Internet of Things (IoT) is a network that links physical items to the Internet and allows them to broadcast data to intelligently identify themselves and enable their location, tracking, monitoring, and management with the assistance of radio frequency identification tags, sensors, actuators, and positioning systems [4]. IoT allows previously inaccessible physical and everyday items to separately gather and share data via the use of built-in sensors and network connection. IoT emerged from networked devices, which started at MIT’s Auto-ID Center in the nineties, according to Center Director Kevin Ashton, who purportedly developed the concept of IoT at the turn of the millennium. The current concept of IoT is supported by the addition of GPS devices, smartphones, social networks, cloud computing, and data analytics to existing networks [5]. IoT plays a significant part in Industry 4.0, which encompasses a variety of innovative and revolutionary technologies, cyber-physical systems, IoT, and cloud computing [6]. It links objects, sensors, actuators, and other smart technologies, allowing immediate access to physical commodity data to support new services with high efficiency and productivity [7]. IoT deployment may boost the likelihood of reliably transferring flow from source to sink, which is the potential of the cold chain network to withstand increased rivalry among enterprises [8].
@@
Improved operating processes, reduced risk and expense, greater visibility and transparency, and greater adaptability and flexibility across the supply chain are merely a few of the many advantages of incorporating IoT into the supply chain [9]. IoT technology may drastically boost the monitoring and control of CSCs, particularly in areas such as medication, food, and logistics where temperature-sensitive commodities are moved and preserved [10].
@@
This study conducts an exhaustive analysis of the barriers that hinder the successful implementation of the IoT in cold supply chains. Integrated interpretive structural modeling and decision-making trial and evaluation laboratory (ISM-DEMATEL) is an extensively used technique for decision-making and analysis in complex systems. ISM is beneficial for building hierarchical links among factors and analyzing the structure of complex systems, whereas DEMATEL delivers a quantitative assessment of the intensity and direction of interactions between factors [11][12]. Both approaches allow for a combination of qualitative and quantitative research, with DEMATEL specifically addressing the feedback loops. They are straightforward and obvious, making them easier to explain and grasp by decision-makers [13][14][15].
@@
The increasing volume of international trade and consumer spending has spurred the rapid growth of supply chains for perishable commodities. The logistics for “cold chain” items, such as perishables and frozen foods, constantly evolve. Global revenue for the sector is forecasted to reach USD 277.69 billion by 2023, growing at a CAGR (cumulative aggregate gross revenue) of 7.24 percent between 2018 and 2023 [16].
@@
A simplified and technology-enhanced logistics method for CSC IoT implementation is illustrated in Fig 1. The figure depicts IoT sensors installed across the CSC to track moisture, temperature, and other important factors, including refrigerated vehicles, warehouses, and storage facilities. The chosen parameters were controlled and monitored using a network of sensors and data-collecting microcontrollers [17]. To provide immediate insight into the state of products in transit, data collected from these sensors are wirelessly connected and relayed in real-time to a centralized system [18]. In sectors such as pharmaceuticals and food, where maintaining the proper temperature is crucial to product quality and safety, IoT devices make sure that temperature-sensitive commodities are stored and transported within a designated temperature range [19][20]. By considering traffic, weather, and supply quality, IoT may help enhance delivery routes [21]. This guarantees that the perishable goods arrive on schedule and in the best possible shape. IoT sensors can monitor inventory levels to minimize overstocking or understocking of temperature-sensitive products, which lowers spoilage and perishable goods losses [22]. Well-organized logistics and transportation systems significantly affect product quality. Adopting a high-quality logistics system, especially for perishable items, involves frequent temperature monitoring, real-time product quality assessment, and shelf life determination [23].
The objectives of the study are to apprehend the following:
@@
    To identify the dominant barriers of IoT implementation in cold supply chains.
@@
    To Investigate the influential relationship among barriers using ISM and DEMATEL technique.
  3. 3.
    To identify the most significant barriers and underlying causes and effects hindering the enhancement of IoT implementation.
Several studies have delved into barriers within specific countries and contexts; however, these investigations are not without their limitations. Some have explored sectors unrelated to the CSC, and while certain studies have focused on IoT adoption in this context, they often fall short of establishing connections among identified barriers or evaluating their respective effects on the overall system. Moreover, the exploration of IoT implementation barriers in CSCs represents a significant gap in the literature, with no prior comprehensive studies identified in this specific domain. The lack of scholarly attention to this crucial aspect of CSC underscores the need for an in-depth investigation. Our research is a pioneering effort to bridge this gap by employing an integrated ISM-DMATEL approach.
The literature gap is particularly evident in the scarcity of empirical examinations and theoretical frameworks dedicated to understanding the intricacies of IoT implementation in CSCs. The unique challenges posed by the CSC environment, characterized by stringent temperature control requirements and perishable goods logistics, warrant further exploration.
To address these gaps, this study contributes to the literature in the following ways:
  1. 1.
    An in-depth analysis of research on essential barriers to CSC IoT adoption, considering various perspectives, including supply chain stages, environmental effects, socio-economic scenery, and data complexity.
  2. 2.
    A survey-based approach was employed to gather data and expert opinions regarding these barriers.
@@
    The MICMAC strategy was applied to comprehensively analyze the fundamental constructs, thereby offering valuable insights into their significance.
@@
    The DEMATEL approach was used to validate causal relationships among the components within the model.
@@
    Key barriers that play a fundamental role in efficiently managing the cold chain are highlighted, offering insights to overcome challenges in IoT adoption.
@@
    Systematic identification and ranking of barriers based on their influence on the system addresses the current lack of comprehensive assessments in the literature.
@@
    Providing a structured approach to understanding and addressing the challenges associated with IoT implementation in CSC, contributing to enhanced operational efficiency and reliability.
@@
    It offers valuable insights for enterprises, governments, and technology developers to navigate and overcome obstacles in integrating IoT into CSC operations.
The rest of the paper is arranged as follows: “Literature review” provides a comprehensive literature review of IoT implementation in CSCs, including implementation challenges and comparable methodologies utilized in the present era; “Methodology” presents the research methodology that was used for the study; “Numerical illustration for results interpretation” covers the data analysis, along with the application of ISM-DEMATEL to the obtained barriers; “Discussions” describes the results of the implementation of ISM-DEMATEL techniques and discusses the results. Finally, the study concludes with conclusions, limitations, and recommendations for future work in “Conclusions and future works.”
@@ 2 Literature review
The aim of this section is to provide a summary of how CSC procedures may use IoT. We used Google Scholar and many other publishers’ websites, including Springer, Elsevier, Emerald, and Taylor-and-Francis, to compile the literature most relevant to our study.
@@ 2.1 Internet of Things (IoT)
The IoT is among the most revolutionary technological developments brought about by the digital era [24]. The Internet of Things is anticipated to become the most popular form of internet traffic in the next years, with a predicted CAGR of 40 in the commercial sector [25]. The economy, communication patterns, and technology development are all expected to be profoundly affected [26]. First invented in 1999 by Kevin Ashton and many other scholars at the Massachusetts Institute of Technology, IoT has gained broad interest because it can provide comfort, efficiency, and competitive benefits to the corporate sector [27]. The IoT is an essential idea in the coming period. Numerous smart buildings and smart houses are built using IoT technologies [28]. The IoT presents a new opportunity to reduce risks, manage complexity, and produce concrete economic benefits by increasing transparency and flexibility across the supply chain [29]. The physical components that link IoT include wireless devices such as portable computers, smartphones, wearable gadgets, RFID tags, RFID readers, and wireless sensors. Users may interact with the IoT network using a variety of devices, including cellphones, implanted RFID transponders, wirelessly connected PCs, and wearable technology [30]. The popularity and success of IoT have skyrocketed in recent years. Industry leaders, technologists, and consumers recognize the importance of IoT, which has led to an industrial revolution and automated operations in organizations and homes [31]. Human-to-thing communication and autonomous coordination among “things” is possible even when they are kept at a facility or transported from one entity in the supply chain to another. Increased supply chain visibility, responsiveness, and adaptability to a variety of SCM challenges are all made possible by IoT [32] [5].
@@ 2.2 IoT in CSC
IoT, often known as the third wave of IT, has revolutionized supply chain management (SCM) and other sectors. This helps expand the supply chain function by offering technical assistance for greater visibility, stability, and intelligent management [33] [34]. Ben-Daya et al. [5] analyzed the existing literature to learn more about the IoT and how it influences SCM. IoT has transformed supply chain management by improving efficiency, allowing predictive maintenance, and enabling real-time reactions to change circumstances, improving operational efficiency, waste reduction, and overall customer happiness [35][35]. Bahl [36] explores how supply chain management has been transformed by IoT technology, which offers real-time data insights, improves visibility, and allows proactive decision-making to satisfy consumer expectations in a changing marketplace and optimize operations. In addition, the majority of research has concentrated on the steps in the supply chain pertaining to food production, distribution, and processing. Kamble et al. [37] proposed an integrated ISM-DEMATEL model to identify the most significant barriers to the widespread deployment of IoT technologies in the food retail supply chain. The two most significant barriers to the widespread use of IoT were a lack of government regulation and insufficient network infrastructure. Nguyen et al. evaluated the reliability of a cold chain network by considering both flow and time factors simultaneously, with the ultimate goal of determining the network reliability for actual products. Shashi [38] revealed two primary themes: the existing cold chain system’s known or unknown limits, as well as the use of digital enablers based on the IoT.
@@
Numerous deployment strategies, constraints, and barriers in cold chain monitoring applications have been discussed [39]. Sallam et al. [34] analyzed future developments in the field of supply chain management and offers a thorough analysis of the IoT applications in this area, including problems, possibilities, and best practices. The feasibility of implementing modern technologies such as IoT to enhance cold SC was assessed by Gupta et al. [40]. To enhance service quality, the organization must invest additional funds for maintenance. Umamaheswari et al. [41] suggested smart cold storage that uses cutting-edge supply chain technology and the IoT to boost productivity and reduce wait times at all levels. Luo et al. [42] presented a smart monitoring system that stores data on defined servers to monitor cold chain items in real time while cutting costs, enhancing network capacity, and simplifying protocols. Cold chain transit is easy to track using this strategy. Afreen and Bajwa [43] implemented a system that utilizes the IoT to track and notify users of changes in cold storage conditions such as humidity, temperature, light, and gas concentration in real-time., alerting workers when these variables exceed safe thresholds. Goodarzian et al. [44] proposed a model for improving the precision, swiftness, and fairness of vaccine injection and implemented an inaugural responsive green-cold vaccine supply chain network in response to the COVID-19 pandemic. Intelligent control technology, cellular networks, wireless communication technology, temperature and humidity sensors, IoT, and database technology all come together to form a cold chain management system[45]. Ceken and Abdurahman [46] presented an IoT-based cold chain logistics system that controls and monitors the cold chain’s ambient temperature in real time and estimates perishable item shelf life to help all stakeholders make better decisions. The paper by Tsang et al. [22] presented an IoT-based risk monitoring solution to handle product quality and safety concerns in the cold chain. Using this technology, the cold chain can be monitored in real-time, and risks can be assessed to ensure the safety of workers. Hernandez and Yamaura [47] suggested that IoT-enabled devices may reduce pharmaceutical waste caused by temperature changes. It also demonstrated consistency in IoT implementation, including infrastructure, security, and device management. Owing to perishable food and cross-regional transit, cold chain systems encounter several challenges in the perishable food supply chains. Atyam et al. explained the complexity and problems associated with modern supply chain management, the importance of IoT in monitoring items, and how IoT and blockchain might improve supply chain performance. Kumar et al. [48] highlighted the need of efficient cost management and waste reduction as key performance indicators and also emphasized the possibility of boosting operational efficiency and optimizing resource utilization via the integration of the IoT. IoT-enabled smart monitoring and control systems could provide long-term solutions to energy consumption and managerial reluctance in the cold supply chain [49]. This study of Zhang et al. [50] addressed these difficulties by creating a perishable food supply chain model utilizing real-time IoT data and two supply hubs to increase performance. Future cold supply networks can become more resilient and sustainable if IoT technologies are used. This advancement is expected to result in less waste and decreased operational expenses, therefore benefiting both businesses and consumers [40] [51].